OpenStudy (anonymous):

1. For the given equation, list the intercepts and test for symmetry. y=x^3−1000. Please show the symmetry test!!

OpenStudy (akashdeepdeb):

y intercept can be found when x = 0 So y intercept = -1000 x intercept can be found when y = 0 So x intercept = ? y=0 = x^3 - 1000 x^3 = 1000 = 10^3 And so x = 10 Understood? :)

OpenStudy (anonymous):

Thanks alot

OpenStudy (akashdeepdeb):

:)

OpenStudy (anonymous):

How do I test for symmetry?